darkknight:

how would you find the max of this: 400/(484t^2-1452t+1189)? and what t is at this max?

justjm:

max would be infinity, right? If you evaluate the vertical asymptote at 1.5, \(\lim_{t\rightarrow\ 1.5}=\infty\)

darkknight:

its 4, though

justjm:

hold on I'm confuzzled it's \(\frac{400}{484x^{2}-1452x+1089}\) right

InsatiableSuffering:

Are you allowed to use the calculator on this? You can simply plug it into y= and then use 2nd trace to find the maximum of the function.

darkknight:

+1189 instead of 1089 but yes

darkknight:

we can use only a scientific calculator and im in precalc not calc

darkknight:

I don't know second trace

darkknight:

no, just not graphing

darkknight:

we can't use a graphing calculator

InsatiableSuffering:

OH WAIT I think I know now

InsatiableSuffering:

You need to use \[\frac{ -b }{ 2a }\] to find the vertex. This allows you topinpoint the ABSOLUTE MINIMUM value of the quadratic, as it opens upwards. Thus, you find your value of t that will satisfy the equation.

InsatiableSuffering:

Then just plug in that value and solve for y

darkknight:

oh lol, time to go back to algebra 1 :)